OpenStudy (fredoneedshelp):

HELP! In the past year, Salma watched 64 movies that she thought were very good. She watched 80 movies over the whole year. Of the movies she watched , what percentage did she think were very good?

OpenStudy (mathmale):

Set up a ratio or proportion: (# "very good") / (total # of movies). Show your work. Lastly, multiply the resulting ratio by 100%. Round off to the nearest %-age point.

OpenStudy (fredoneedshelp):

@mathmale is it 0.8%?

OpenStudy (mathmale):

Please show the ratio first: 64/80. Next, multiply your result by 100%. 0.8% is a very small percentage and thus could not be right. This viewer considered most of the movies to be "very good." find 64/80, and then multiply the result by 100%.

OpenStudy (fredoneedshelp):

OHHH okAY

OpenStudy (fredoneedshelp):

64/80 is 0.8 @mathmale

OpenStudy (mathmale):

correct. What do you need to do next? (Hint: I've suggested the required next step twice.)

OpenStudy (fredoneedshelp):

i did 0.8*100% and still got 0.8

OpenStudy (fredoneedshelp):

@mathmale

OpenStudy (wilsondanielle):

don't multiply by 100% (which is equal 1) - multiply by 100. This converts it to a percentage (i.e. how many out of 100). It's like cross multiplication - 64/80 = x/100

OpenStudy (mathmale):

The original problem statement asks that you find "the percentage of the 80 movies that this viewer watched that he / she rated as very good." Technically, the answer is a percentage, and requires the % symbol. You need to convert 64/80 to a percentage. The instructions I have given you are correct. 64/80 = 0.8: correct intermediate answer (64/80)*100% = 80% correct final answer

OpenStudy (fredoneedshelp):

okay i see

OpenStudy (fredoneedshelp):

i had that answer i wasnt sure if it was right

OpenStudy (mathmale):

wilsondanielle Best Response Medals 0 don't multiply by 100% (which is equal 1) - multiply by 100. This converts it to a percentage (i.e. how many out of 100). Danielle: I disagree with you. Before you contradict another person trying to help here, please make sure you have looked up the content to ensure that you have not misunderstood the situation. Thank you.

OpenStudy (mathmale):

If not asked for a "percentage" answer, you could respond: 1) four fifths of the films were rated "very good" or 80% of the films were rated "very good."